How many kgs of salt must be added to 24 kg of 20% salt solution in order to increase the concentration of salt to 40%?
Answer by mananth(16946)

How many kgs of salt must be added to 24 kg of 20% salt solution in order to increase the concentration of salt to 40%?

salt 100%-------------------- 20%---------------40% strength
x------------------------------24 kgs-----------x+24 kgs
100%*x+20%*24=40%*(x+24)
1x+4.80=.4x+9.6
x-0.4x=6-4.8
0.6x=1.2
x=1.2/0.6 = 2 kgs